OpenStudy (zaarin):

On a blueprint, 1/4 inch represents 1 foot. If a window is supposed to be 60 inches wide, how wide would its representation be on the blueprint?

OpenStudy (anonymous):

The ratio remains the same, so:\[ \frac{1}{1/4}=\frac{x}{60} \]

OpenStudy (anonymous):

Cross multiply: \[ \frac 14 x = 60 \]

OpenStudy (zaarin):

sorry that's not the answer.

OpenStudy (anonymous):

I didn't say it was. You need to solve for \(x\).

OpenStudy (zaarin):

exaclty the answer is supposed to be 5/4 but how I don't know.